Suppose the data below represents the entire production of an economy, where the table shows the
production of goods and their prices for 2012 and 2013:
(a) Calculate nominal GDP for each year. Calculate the percentage growth rate of nominal GDP.
(b) Calculate real GDP for each year using 2012 as the base year. Calculate its growth rate.
(c) Calculate the GDP deflator for each year and its inflation rater. Briefly explain the relationship between the inflation rate and the growth rates of nominal and real GDP.
Please HELP! I am struggling with this!!

Good 1 Good 2
2012 {quantity, price/unit} {400, $1.00} {200, $2.00}
2013 {quantity, price/unit} {420, $0.90} {215, $2.20}
1
Expert's answer
2015-02-24T09:26:44-0500
 Good 1 Good 2
2012 {400, $1.00} {200, $2.00}
2013 {420, $0.90} {215, $2.20}
(a) Nominal GDP for each year is:
GDP(2012) = 1*400 + 2*200 = $800
GDP(2013) = 0.9*420 + 2.2*215 = $851
The percentage growth rate of nominal GDP is (851/800 - 1)*100% = 6.38%
(b) Real GDP(2012) = $800
Real GDP(2013) = 1*420 + 2*215 = $850
GDP growth rate = (850/800 - 1)*100% = 6.25%
(c) The GDP deflator for 2012 is 1 and for 2013 is 851/850 = 1.0012
So, the inflation rate is 0.12%.